9. As Matt’s consumption of Santa Cruz (sparkling water) goes up, his ____.

A) average utility from consuming Santa Cruz increases.

B) total utility from consuming Santa Cruz increases.

C) marginal utility from consuming Santa Cruz increases.

D) elasticity of utility from consuming Santa Cruz increases.

10. Diminishing marginal utility means that ____.

A) Taylor will enjoy his second taco less than the first.

B) the total utility from one taco exceeds the total utility from two hamburgers.

C) the price of two taco s is twice the price of one.

D) beyond a certain point, total utility decreases as income rises.

11. Jessica likes to buy only two goods: Cheetos and Pepsi. To determine how Jessica can maximize her utility from consuming Cheetos and Starbucks Frappuccino, you need to know

I. Jessica’s preferences for Cheetos and Frappuccino.

II. The prices of Cheetos and Frappuccino.

III. Jessica’s income.

A) I only.       B) I and II.        C) II and III.        D) I, II and III.

12. Lauren likes the combination of Larry's Pizza and Dr. Pepper (soda) for dinner. When the price of a soda is $2 and the price of a slice of pizza is $4, Lauren maximizes her utility by buying 5 sodas and 10 slices of pizza. If the marginal utility of the 5th soda is 100 units of utility, then the marginal utility of the 10th slice of pizza must be ____ units of utility.

A) 200

B) 100

C) 50

D) More information is needed to determine the marginal utility of the 10th slice of pizza.
Answer #1

9. B) total utility from consuming Santa Cruz increases.
(With increase in consumption, TU will increase.)

10. A) Taylor will enjoy his second taco less than the first.
(Diminishing MU means that the utility from second unit will be less than the utility from the first unit.)

11. D) I, II and III
(All three information are required to maximize the utility.)

12. A) 200
(Utility is maximized when MU/P for two goods is equal.
So, 100/2 = MU of pizza/4
So, MU of pizza = (100/2)*4 = 200)